A Japanese Soft Drink Company is planning to establish a subsidiary company in India to produce Mineral Water. Based on the estimated annual sales of 40,000 bottles of the mineral water, cost studies produced the following estimates for the Indian subsidiary. % of Variable cost Material Rs. 1,20,000 100% Labour Rs. 1,50,000 80% Factory Overhead Rs. 92,000 60% Administration O/H Rs. 40,000 35% The Indian production will be sold by manufacturer‟s representatives who will receive a commission of 8% of the sale price. No portion of the Japanese office is to be allocated to the Indian Subsidiary. Required to a) Compute the sale price per bottle enable the management to realise an estimated 10% profit on sale proceeds in India. b) Calculate the BEP in rupees and also in number of bottles for the Indian subsidiary on the assumption that the sale price is Rs. 14/- per bottle. c) Prepare break even chart and comment on the profitability of the firm. d) Discuss the relevance of CVP analysis in this context.
Answer #1

Note: It is assumed that the costs given in the question are Total Costs per annum and the % given represents the variable portion and remaining the fixed portion.

Particulars Total Annual Cost Variable Cost % Variable Cost Fixed Cost
Materials 120000 100 120000 -
Labour 150000 80 120000 30000
Factory Overhead 92000 60 55200 36800
Admin Overhead 40000 35 14000 26000
TOTAL 402000 309200 92800

a) Computation of Selling Price per unit of bottle to realise a profit of 10% on Sales

Let the Selling Price per unit be x.

Annual Sales = 40000x

Commission (8% on Sales) = 3200x

Target Profit (10% on Sales) = 4000x

Total Cost (from above table) = 402000

Sales = Cost + Profit

40000x = 402000+3200x+4000x

Solving, x = 12.256

Thus, Selling Price per unit of bottle to realise a profit of 10% on Sales = 12.256 per unit

b) Required Break-even point in Sales and Quantity:

Assumption, Sales Price per bottle = Rs. 14

BEP in Sales = Total Fixed Costs/Contribution Sales Ratio(PV Ratio)

BEP in Quantity = Total Fixed Costs/Contribution per unit

  • Total Fixed Costs = 92800
  • Contribution per unit = Selling Price per unit - Variable Cost per unit

Variable Cost per unit = 309200/40000 = 7.73

Sales Commission per unit = 14*8% = 1.12

Contribution = 14-7.73-1.12 = 5.15 per unit

  • Contribution Sales Ratio = 5.15/14 = 36.79%

BEP (Sales) = 92800/36.79% = Rs. 252242

BEP (Quantity) = 92800/5.15 = 18019 units

d) Relevance of CVP Analysis:

  • CVP analysis provides relationship between cost, volume and profit and hence becomes important for budgeting and planning.
  • It helps management to identify most profitable combination of costs and volume.
  • We can also identify how change in costs and volume effect changes in income.
  • Hence it plays a major role in decision making.

PS: I am not sure about how to draw a Breakeven Chart (Part (c) of the question)

Thanks.